A landscaper needs to mix 75% pesticide solution with 30 gals. Of 25% pesticide solution to obtain a 60% pesticide solution. How many gallons of the 75% solution must be used ?
Answer by lwsshak3(11628)

A landscaper needs to mix 75% pesticide solution with 30 gals. Of 25% pesticide solution to obtain a 60% pesticide solution. How many gallons of the 75% solution must be used
***
let x=gal of 75% solution must be used
x+30=gal of 65% solution
..
75%*x+25%*30=60%*(x+30)
.75x+.25*30=.6x+.6*30
.75x+7.5=.6x+18
.15x=10.5
x=10.5/.15=70
ans:
70 gal of 75% solution must be used